If x varies directly as 3y and x=8 when y=3 find x when y=27
Answer by ewatrrr(24785)

x varies directly as 3y
x = k(3y)
x=8 when y=3
8 = k(9)
8/9 = k
find x when y=27
x = (8/9)(3y)
x = (8/9) (3*27)
x= 8*9 = 72